OpenStudy (anonymous):

f(x) = x^2 [x] then f`(3/2) = ?

OpenStudy (fibonaccichick666):

what is the derivative of \(x^2 x\)?

OpenStudy (fibonaccichick666):

or better yet, can you simplify \(x^2x\)?

OpenStudy (anonymous):

I think [x] is supposed to be the floor function

OpenStudy (fibonaccichick666):

oh uhm then I am no help unfortunately... I don't even think that has a derivative does it? It's no where near continuous

OpenStudy (anonymous):

since [x] = 1 when 1 <= x < 2 x^2 [x] = x^2, for 1 <= x < 2 so, f'(x) = 2x = 2(3/2) = 3

OpenStudy (fibonaccichick666):

ahhh if you define it piecewise it does

OpenStudy (anonymous):

yes :)

OpenStudy (fibonaccichick666):

Idk I guess we shall see when @hamoody1996 responds

OpenStudy (primeralph):

[x] should be the int function?

OpenStudy (freckles):

so f'(x)=floor(x)*2x when x is not equal to an integer i believe

OpenStudy (fibonaccichick666):

but anyways we don't know if it's the floor since they have not responded, and [] are not how you represent floor, floor won't have the toppart soo..... waiting on a response......

OpenStudy (freckles):

it looks like it can mean either http://mathworld.wolfram.com/NearestIntegerFunction.html

OpenStudy (freckles):

depending on what the person who wrote it meant

OpenStudy (fibonaccichick666):

floor=int

OpenStudy (fibonaccichick666):

sort of

OpenStudy (freckles):

no i was talking about the notation

OpenStudy (freckles):

it says in that link [ ] could mean int or floor

OpenStudy (freckles):

i wasn't saying they are equal

OpenStudy (fibonaccichick666):

oh

OpenStudy (anonymous):

I believe the convention is that [x] is the floor function, not the ceiling function.

OpenStudy (fibonaccichick666):

we still don't know if that is even what they were asking haha

OpenStudy (fibonaccichick666):

they're offline, guess we shall never know

OpenStudy (primeralph):

If anything, it should be the floor function or integer part.

OpenStudy (anonymous):

as far as I know, [x] = int[x] = floor(x).

OpenStudy (anonymous):

actually [x] is the floor function and the choices were 1 ,2 , 3 , not found thanks yall